# A Continual Offline Reinforcement Learning Benchmark for Navigation Tasks

### v1

@@ Abstract
Index Terms:
Deep Learning, Offline Reinforcement Learning, Continual Learning, Bots, Benchmarks, Human Generated Data.
@@ III-C Continual Reinforcement Learning
In CRL, an agent follows a sequence of tasks, or stream, $\mathcal{T}=\big{(}T_{1},...,T_{N}\big{)}$, with $T_{k}=(\mathcal{M}_{k},\mathcal{D}_{k})$. We note as $\theta_{k}$ the parameters of a policy after learning on the $k$-th task. As the agent learns new skills, it must preserve (to prevent forgetting) and enhance (to encourage backward transfer) its performance on tasks already learned, while ideally having a relatively low number of parameters. To quantitatively compare CRL methods, we adopt standard metrics commonly used in the literature [15, 26] : Performance (PER) : $\frac{1}{N}\sum_{k=1}^{N}\sigma_{\mathcal{M}_{k}}(\theta_{N})$; Backward Transfer (BWT) : $\frac{1}{N}\sum_{k=1}^{N}\Big{(}\sigma_{\mathcal{M}_{k}}(\theta_{N})-\sigma_{\mathcal{M}_{k}}(\theta_{k})\Big{)}$; Forward Transfer (FWT) : $\frac{1}{N}\sum_{k=1}^{N}\Big{(}\sigma_{\mathcal{M}_{k}}(\theta_{k})-\sigma_{\mathcal{M}_{k}}(\tilde{\theta}_{k})\Big{)}$ ; Relative Model Size (MEM) : $\frac{|\theta_{N}|}{|\theta_{\text{ref}}|}$; Model Inference Cost (INF); Model Training Cost (TRN). PER measures the success rate across all tasks. BWT indicates how learning a new task affects previous ones, while FWT measures the transfer of knowledge, using $\tilde{\theta}_{k}$ as randomly initialized parameters. MEM compares the memory load of the model to a reference one associated to parameters $\theta_{\text{ref}}$. INF is the duration required for policy inference across tasks, and TRN is the overall training duration.
@@ IV-A Navigation Environments & Mazes
We propose video game–inspired navigation environments, which offer easy to use 3D maze configurations via our provided open-source code [^1] GitHub Repository Link.. Our benchmark comprises two families of mazes designed to evaluate CRL agents : SimpleTown, a collection of $8$ relatively simple mazes ($20\ m\times 20\ m$), and AmazeVille, a more complex set of $8$ mazes ($60\ m\times 60\ m$).Table IV-A and Table IV-A display the available outputs and inputs.
{tblr}
colspec = |c||c|c|c|, row1 = SeaGreen,
@@
Move Forwards   $1$   bool  Z
  Move Backwards   $1$   bool  S
  Move Left   $1$   bool  Q
  Move Right   $1$   bool  D
  Jump   $1$   bool  Space
  Turn   $1$   float  Mouse
@@
Position $3$ float Orientation $3$ float
Goal Position $3$ float Velocity $3$ float
@@
Floor Contact $1$ bool Wall Contact $1$ bool
Goal Contact $1$ bool Timestep $1$ int
Up Direction $3$ float - - -
TABLE II: Available observation features. Position data are in $m$, while the Orientation is in $rad$, and the Velocity in $m\cdot s^{-1}$. The RGB Images offer a visualization of the agent’s viewpoint, and the Depth Image is generated from $11\times 1$ 11 raycasts thrown towards nearest obstacles. Although our experiments only use positions and depth images, we also provide the pixel-based observations with all datasets.
@@ V-C Implementation Details
We base our configuration on the approach described in [29]. In all our experiments, we use Residual MLPs [30] with Layer Normalization [31] for the hidden layers. All networks use3×25632563$\times$ 2563 × 256 hidden units, GELU activations, and are initialized with a variance scaling strategy [32] (with a scale of $0.1$).
@@ VI-A Offline Reinforcement Learning
{tblr}
colspec=|c||c|c||c|c|, row1 = Melon,
[…]
@@
A – HOOX 73.2 $\pm$
7.5 78.1 $\pm$ 5.2 222.3 $\pm$ 7.8 215.0 $\pm$ 9.7
A – HXOO 92.7 $\pm$
2.1 97.9 $\pm$ 2.3 185.7 $\pm$ 5.3 181.4 $\pm$ 5.0
A – HXOX 79.4 $\pm$
2.8 85.4 $\pm$ 3.4 239.4 $\pm$ 2.0 234.2 $\pm$ 3.1
A – LOOO 92.7 $\pm$ 3.8 88.5 $\pm$
3.0 138.0 $\pm$ 5.0 142.0 $\pm$ 4.5
A – LOOX 85.5 $\pm$ 7.9 82.7 $\pm$
5.7 171.8 $\pm$ 12.6 173.0 $\pm$ 9.7
A – LXOO 89.2 $\pm$
3.5 90.1 $\pm$ 3.8 145.2 $\pm$ 6.0 141.3 $\pm$ 7.2
A – LXOX 93.4 $\pm$ 2.2 93.4 $\pm$ 2.2 160.4 $\pm$ 3.8 161.8 $\pm$
3.9
@@ VI-B1 Performance
Table VI-B1 presents the performance metric of each CRL method on the task streams. To a certain extent, PER simultaneously tells about generalization and forgetting.
{tblr}
colspec=|c||c|c||c|c||c|c|, row1 = Melon,
@@ VI-B2 Backward & Forward Transfer
Figure 5: Backward and Forward Transfer Metrics. Architectural or separate-policy methods (PNN, HiSPO, SCN, FTN) typically balance old-task retention and new-task adaptation, whereas single-policy or regularized methods (SC1, FRZ, EWC, L2) risk higher forgetting or reduced forward transfer in complex navigation scenarios.
@@ VI-B3 Relative Model Size
These results highlight a core trade-off : methods like PNN or SCN & FTN can improve continual learning but incur greater memory overhead, while all the single-model approaches (such as SC1, FRZ, RPL) remain lightweight but risk higher forgetting. Balancing memory cost and adaptation is thus key to deploying continual offline reinforcement learning in practical pipelines.
@@ VI-B4 Training & Inference Costs
Figure 8: Average INF (model inference cost duration in ms). We measure inference time averaging on 100,000 forward passes on CPU with a batch size of 64, using the AR1 stream (seed 100).At inference, single-model approaches (SC1-N, FT1-N, FRZ, RPL, EWC, and L2) incur minimal overhead, requiring a single forward pass per step. In contrast architectural strategies (PNN and HiSPO) may demand additional computations (such as selecting the right sub-policy or combining anchors), leading to higher inference times.
Figure 9: Average TRN (training duration in min). Single-model approaches train quickly by updating a single parameter set, while RPL is slightly slower due to repeated dataset sampling. EWC and L2 add overhead by storing parameters snapshots. Architectural methods also take longer to train, as PNN shows non-linear growth by adding new columns and lateral connections per task, and HiSPO introduces new anchors and also computes cosine-similarity regularization weights.
Overall, these findings show that single-model methods maintain lighter overhead during both training and inference but can suffer from higher forgetting. Conversely, multi-policy or subspace-based strategies better preserve knowledge at the cost of increased computational demands, a crucial consideration in large-scale or time-sensitive applications.
@@ VII Discussion & Future Work
We introduced Continual NavBench, a new benchmark designed for Continual Offline Reinforcement Learning in video game–inspired navigation tasks. Our experiments show that while approaches like PNN and HiSPO are good at preserving and reusing skills across tasks, they may incur either significant memory or computational costs. In contrast, single-model or regularized methods (SC1, FRZ, EWC, L2) remain lighter in resource consumption but often exhibit higher forgetting. These findings highlight the inherent trade-offs between scalability, performance, and efficiency in Offline CRL.
Looking ahead, we believe there are several promising paths for future research. First, enlarging the environment suite (e.g., multi-level mazes, dynamic obstacles, or different agents) could further stress the agent’s ability to adapt in complex scenarios. Second, exploring hybrid approaches (e.g., combining subspace methods with replay or regularization) may yield more balanced solutions that preserve prior knowledge without increasing memory or training costs. Third, incorporating domain transfer, from offline learning to online finetuning, to shed light on how efficiently each method generalizes beyond using datasets.
Ultimately, we hope Continual NavBench spurs more rigorous and reproducible evaluations of continual learning strategies for navigation and beyond. By offering diverse tasks, open-source tools, and well-defined metrics, our benchmark lowers the barrier for future work to compare methods, refine architectures, and propose novel solutions. We believe that systematically addressing memory, training, and inference constraints will be pivotal in bringing lifelong autonomy to both realistic research prototypes and real-world production pipelines.

### v2

@@ III-C Continual Reinforcement Learning
In CRL, an agent follows a sequence of tasks, or stream, $\mathcal{T}=\big(T_{1},...,T_{N}\big)$, with $T_{k}=(\mathcal{M}_{k},\mathcal{D}_{k})$. We note as $\theta_{k}$ the parameters of a policy after learning on the $k$-th task. As the agent learns new skills, it must preserve (to prevent forgetting) and enhance (to encourage backward transfer) its performance on tasks already learned, while ideally having a relatively low number of parameters. To quantitatively compare CRL methods, we adopt standard metrics commonly used in the literature [5, 12] : Performance (PER) : $\frac{1}{N}\sum_{k=1}^{N}\sigma_{\mathcal{M}_{k}}(\theta_{N})$; Backward Transfer (BWT) : $\frac{1}{N}\sum_{k=1}^{N}\Big(\sigma_{\mathcal{M}_{k}}(\theta_{N})-\sigma_{\mathcal{M}_{k}}(\theta_{k})\Big)$; Forward Transfer (FWT) : $\frac{1}{N}\sum_{k=1}^{N}\Big(\sigma_{\mathcal{M}_{k}}(\theta_{k})-\sigma_{\mathcal{M}_{k}}(\tilde{\theta}_{k})\Big)$ ; Relative Model Size (MEM) : $\frac{|\theta_{N}|}{|\theta_{\text{ref}}|}$;
Model Inference Cost (INF); Model Training Cost (TRN). PER measures the success rate across all tasks. BWT indicates how learning a new task affects previous ones, while FWT measures the transfer of knowledge, using $\tilde{\theta}_{k}$ as randomly initialized parameters. MEM compares the memory load of the model to a reference one associated to parameters $\theta_{\text{ref}}$. INF is the duration required for policy inference across tasks, and TRN is the overall training duration.
@@ IV-A Navigation Environments & Mazes
We propose video game–inspired navigation environments, which offer easy to use 3D maze configurations via our provided open-source code [^1] GitHub Repository Link.. Our benchmark comprises two families of mazes designed to evaluate CRL agents : SimpleTown, a collection of $8$ relatively simple mazes ($20\ m\times 20\ m$), and AmazeVille, a more complex set of $8$ mazes ($60\ m\times 60\ m$).Table I and Table II display the available outputs and inputs.
@@
Move Forwards 11 bool Z
Move Backwards 11 bool S
Move Left 11 bool Q
Move Right 11 bool D
Jump 11 bool Space
Turn 11 float Mouse
@@
Position 33 float Orientation 33 float
Goal Position 33 float Velocity 33 float
@@
Floor Contact 11 bool Wall Contact 11 bool
Goal Contact 11 bool Timestep 11 int
Up Direction 33 float - - -
TABLE II: Available observation features. Position data are in $m$, while the Orientation is in radrad, and the Velocity in $m\cdot s^{-1}$. The RGB Images offer a visualization of the agent’s viewpoint, and the Depth Image is generated from $11\times 11$ raycasts thrown towards nearest obstacles. Although our experiments only use positions and depth images, we also provide the pixel-based observations with all datasets.
@@ V-C Implementation Details
[…]
@@ VI-A Offline Reinforcement Learning
Task Success Rate Episode Length
@@
A – HOOX 73.2 $\pm$ 7.5 78.1 $\pm$ 5.2 222.3 $\pm$ 7.8 215.0 $\pm$ 9.7
A – HXOO 92.7 $\pm$ 2.1 97.9 $\pm$ 2.3 185.7 $\pm$ 5.3 181.4 $\pm$ 5.0
A – HXOX 79.4 $\pm$ 2.8 85.4 $\pm$ 3.4 239.4 $\pm$ 2.0 234.2 $\pm$ 3.1
A – LOOO 92.7 $\pm$ 3.8 88.5 $\pm$ 3.0 138.0 $\pm$ 5.0 142.0 $\pm$ 4.5
A – LOOX 85.5 $\pm$ 7.9 82.7 $\pm$ 5.7 171.8 $\pm$ 12.6 173.0 $\pm$ 9.7
A – LXOO 89.2 $\pm$ 3.5 90.1 $\pm$ 3.8 145.2 $\pm$ 6.0 141.3 $\pm$ 7.2
A – LXOX 93.4 $\pm$ 2.2 93.4 $\pm$ 2.2 160.4 $\pm$ 3.8 161.8 $\pm$ 3.9
@@ VI-B1 Performance
Table IV presents the performance metric of each CRL method on the task streams. To a certain extent, PER simultaneously tells about generalization and forgetting.
@@ VI-B2 Backward & Forward Transfer
Figure 5: Backward and Forward Transfer Metrics. Architectural or separate‐policy methods (PNN, HiSPO, SCN, FTN) typically balance old‐task retention and new‐task adaptation, whereas single‐policy or regularized methods (SC1, FRZ, EWC, L2) risk higher forgetting or reduced forward transfer in complex navigation scenarios.
@@ VI-B3 Relative Model Size
These results highlight a core trade‐off : methods like PNN or SCN & FTN can improve continual learning but incur greater memory overhead, while all the single‐model approaches (such as SC1, FRZ, RPL) remain lightweight but risk higher forgetting. Balancing memory cost and adaptation is thus key to deploying continual offline reinforcement learning in practical pipelines.
@@ VI-B4 Training & Inference Costs
Figure 8: Average INF (model inference cost duration in ms). We measure inference time averaging on 100,000 forward passes on CPU with a batch size of 64, using the AR1 stream (seed 100).At inference, single‐model approaches (SC1-N, FT1-N, FRZ, RPL, EWC, and L2) incur minimal overhead, requiring a single forward pass per step. In contrast architectural strategies (PNN and HiSPO) may demand additional computations (such as selecting the right sub‐policy or combining anchors), leading to higher inference times.
Figure 9: Average TRN (training duration in min). Single-model approaches train quickly by updating a single parameter set, while RPL is slightly slower due to repeated dataset sampling. EWC and L2 add overhead by storing parameters snapshots. Architectural methods also take longer to train, as PNN shows non-linear growth by adding new columns and lateral connections per task, and HiSPO introduces new anchors and also computes cosine‐similarity regularization weights.
Overall, these findings show that single‐model methods maintain lighter overhead during both training and inference but can suffer from higher forgetting. Conversely, multi‐policy or subspace‐based strategies better preserve knowledge at the cost of increased computational demands, a crucial consideration in large‐scale or time‐sensitive applications.
@@ VII Discussion & Future Work
We introduced Continual NavBench, a new benchmark designed for Continual Offline Reinforcement Learning in video game–inspired navigation tasks. Our experiments show that while approaches like PNN and HiSPO are good at preserving and reusing skills across tasks, they may incur either significant memory or computational costs. In contrast, single‐model or regularized methods (SC1, FRZ, EWC, L2) remain lighter in resource consumption but often exhibit higher forgetting. These findings highlight the inherent trade‐offs between scalability, performance, and efficiency in Offline CRL.
Looking ahead, we believe there are several promising paths for future research. First, enlarging the environment suite (e.g., multi‐level mazes, dynamic obstacles, or different agents) could further stress the agent’s ability to adapt in complex scenarios. Second, exploring hybrid approaches (e.g., combining subspace methods with replay or regularization) may yield more balanced solutions that preserve prior knowledge without increasing memory or training costs. Third, incorporating domain transfer, from offline learning to online finetuning, to shed light on how efficiently each method generalizes beyond using datasets.
Ultimately, we hope Continual NavBench spurs more rigorous and reproducible evaluations of continual learning strategies for navigation and beyond. By offering diverse tasks, open‐source tools, and well‐defined metrics, our benchmark lowers the barrier for future work to compare methods, refine architectures, and propose novel solutions. We believe that systematically addressing memory, training, and inference constraints will be pivotal in bringing lifelong autonomy to both realistic research prototypes and real‐world production pipelines.
VIII Acknowledgement
This project was provided with computer and storage resources by GENCI at IDRIS thanks to the grant 2024-AD011015210 on the supercomputer Jean Zay’s CSL partition.